OpenStudy (anonymous):

Solve the following system of equations. y = x^2 + 3 y = x + 5

OpenStudy (anonymous):

I know that the coeffecients must be opposite..

OpenStudy (lυἶცἶ0210):

Since y =x+5 you could plug that into the first equation: \(\large y=\color{green}{x^2+3} \) \(\large y=\color{red}{x+5}\) \(\Large \color{red}{x+5}=\color{green}{x^2+3}\) Now solve

OpenStudy (campbell_st):

so you are looking for the points of intersection, so you can equate them \[x + 5 = x^2 + 3\] if you rewrite the equation you get \[x^2 - x - 2 = 0\] this can be solved by factoring

OpenStudy (anonymous):

okay, would it be 1 and 1? not sure

OpenStudy (campbell_st):

so find the factors of -2 that add to -1 the larger factor is negative and the smaller is positive

OpenStudy (campbell_st):

when you get the solutions for x, substitute them into either equation to find the corresponding y values... remember the solutions are points (x,y)